OpenStudy (anonymous):

PLZ PLZ, SOMEONE HELP ME ON THIS PROBLEM!!! Sally finklestein has a recipe that calls for horseradish sauce that is 45% pure. At the grocery store she finds one horseradish sauce that is 30% pure and another one that was 80% pure. How many teaspoons of each of these sauces should she mix together to get 4 teaspoons of horseradish sauce that is 45% pure horseradish?

OpenStudy (anonymous):

let a = purity of sauce a = 0.3 let b = purity of sauce b = 0.8 x = number of spoonfuls of a y = number of spoonfuls of b x0.3 * y0.8 = 0.45 Total number of spoonfuls = 4 x+y=4 x=4-y Sub 4-y(0.3) * y0.8 = 0.45 Solve for y Sub to solve for x

OpenStudy (anonymous):

Woops the equation i think is supposed to be x0.3 + y0.8 = 0.45 4-y(0.3) + y0.8 = 0.45

OpenStudy (anonymous):

thank you